OpenStudy (anonymous):

particular brand of gasoline has a density of 0.737 g/mL at 25 ∘C. How many grams of this gasoline would fill a 14.6 gal tank (1US gal=3.78L)?

OpenStudy (hlilly2413):

Forgive me if I am incorrect, but I believe you would use the formula for density. \[Density = \left(\begin{matrix}\frac{ mass (g) }{ volume (mL) }\\ \end{matrix}\right)\] So, step 1 would be to convert your 3.78L into mL. To do that you can either just move your decimal 3 places to the right or use: \[3.78L \times \left(\begin{matrix}\frac{ 1000 mL }{ 1L} \\ \end{matrix}\right)\] Liters cancel out and you are left with mL. Step 2, would be to plug your values for density (which was given) and your value for volume (which you figured out in step 1) into the density equation and use algebra to solve for mass. Ex: \[0.373 g/mL = \frac{ x }{ volume (mL )from step 1 }\] To help with the algebra: multiply your density by the volume, mL will cancel out and your answer will be how many grams fit into that volume. Hope this helps!

OpenStudy (hlilly2413):

Value for density*